OpenStudy (anonymous):

find the quotient. 48a^3 bc^2 ÷ 3abc

OpenStudy (aum):

\(\Large \frac{48a^3 bc^2}{3abc} = \frac{48}{3}*\frac{a^3}{a}*\frac{b}{b}*\frac{c^2}{c} = ?\)